# Exploring strong electronic correlations in the breathing kagome metal Fe3Sn

### v1

@@ III Magnetic properties
The amplitude of the DM vector as a function of the interatomic distance is shown in Figure 2(b), while the components Dx, Dy, and Dz for selected exchange paths are given in Table 2. The largest coupling is interlayer, connecting Fe atoms along the path E2, and amounts to 0.52 meV, mostly arising from its Dz component. Inside the kagome plane, the first and second shells of NNs are characterized by slightly smaller DM vectors, whose major contributions still come from the Dz component. These finite values of the DM interaction indicate the tendency for chirality between two Fe sites with the preferred direction for the DM vector along the z-axis (out-of-plane). The magnitude of the DM interaction is relatively small (3-4%) compared to the isotropic Heisenberg exchange tensor, making the likelihood of spin canting quite low Camley and Livesey (2023). Furthermore, we see that |D| for the forth and fifth NN shells is zero, because of the presence of inversion symmetry, as evidenced by the interlayer exchange paths E4 and E5 depicted in Figure 1(c).
| Exchange path | DM vector | DFT | DFT+DMFT | DFT+U |
| E1 | Dx | 0 | 0 | 0 |
| Dy | 0 | 0 | 0 | |
| Dz | 0.41 | 0.22 | 0.50 | |
| E2 | Dx | -0.08 | -0.05 | -0.05 |
| Dy | 0.04 | 0.02 | 0.02 | |
| Dz | 0.51 | 0.19 | 0.36 | |
| E3 | Dx | 0 | 0 | 0 |
| Dy | 0 | 0 | 0 | |
| Dz | 0.28 | 0.56 | 0.78 | |
| E6 | Dx | -0.16 | -0.27 | -0.31 |
| Dy | 0 | 0 | 0 | |
| Dz | 0 | 0 | 0 | |
| E9 | Dx | 0.09 | 0.09 | 0.13 |
| Dy | 0.16 | 0.17 | 0.23 | |
@@
The inclusion of an explicit Hubbard term via DFT+U and DFT+DMFT leads to a small reduction of the isotropic exchange, as shown in Figure 2(a). This reduction reflects the smaller values of the Fe spin moments reported in Table 1 through the definition used in Eqn. 2. By inspecting the projected density of states (PDOS) for the Fe- states, depicted in Figure 3, we observe that the exchange splitting is largest in DFT and smallest in DFT+DMFT, in agreement with the trends of the magnetic moments and interatomic interactions. The overall reduction of the strength of the magnetic coupling in DFT+DMFT is similar to the one previously observed for bcc Fe Chadov et al. (2008); Kvashnin et al. (2015). Nevertheless, it is surprising that the DFT+U does not cause a stronger localization of the Fe- electrons, making the ’s shorter ranged. Although a possible reason for the observed behavior may be the usage of smaller Coulomb interaction parameters in DFT+U, additional calculations demonstrate that this is not the leading factor (data not shown). In fact, the long-range nature is found to be a consequence of the AMF double counting, which is usually used for correlated metals Petukhov et al. (2003). To evaluate the impact of the double counting, we performed additional electronic structure calculations in DFT+U in the fully-localized limit (FLL) while keeping the other settings unchanged. As expected, a much stronger electronic localization is observed, resulting in an increase of the exchange splitting (data not shown), a higher local magnetic moment of 2.54 B, and an increase of the first NN coupling up to 31 meV. This magnetic moment deviates significantly from the experimentally reported values, which demonstrates that the FLL double counting is not suitable for describing this system, as usual for itinerant ferromagnets Petukhov et al. (2003).
@@
Looking at the DM interaction, we witness significant changes, qualitatively and quantitatively, with the inclusion of the on-site Coulomb interaction via DFT+DMFT and DFT+U. For both methods, Figure 2(b) shows a marked increase in the value of |D| for the third shell of NNs, corresponding to the intralayer exchange path E3. The first shell of NNs, along the intralayer exchange path E1, exhibits a more complex trend, as the DM interaction increases in DFT+U but decreases in DFT+DMFT. The dominant character of the intralayer terms remains Dz, compatibly with the symmetry of the kagome lattice. The DM interaction with the second shell of NNs, along the exchange path E2, decreases for both DFT+U and DFT+DMFT, an effect driven mainly by a change in Dz. Notably, there is also a significant increase in the magnitude of the DM interaction with farther NNs, as e.g. in the sixth and ninth shells, which we attribute to a stronger orbital polarization driven by the inclusion of the on-site Coulomb interaction. These data, and in particular the decomposition shown in Table 2, suggest deeper changes in the electronic structure than those observed from the PDOS of Figure 3. We will investigate these changes in the next section.
@@
For completeness, we also calculated the values of the interatomic exchange interactions for the non-relativistic case, corresponding to Eqn. 1. As illustrated in Figure S1 of the SM sup , the isotropic exchange is only marginally affected by the presence of SOC, and the primary consequence of the latter is to induce anisotropic terms as the DM interaction, favoring spin canting and complex magnetic textures Karmakar et al. (2023).
@@ IV Details of the electronic structure
The spectral functions along high-symmetry directions in the Brillouin zone are reported in Figure 4 for DFT, DFT+DMFT and DFT+U, with and without the inclusion of SOC. As in the previous Section, the data for SOC are shown for a magnetization along the easy axis (100). Focusing on the DFT results, panels (a) and (d), we see the presence of several bands with a relatively flat dispersion near the Fermi level. For convenience, we highlighted those bands in magenta and blue colors, reflecting their spin character. In agreement with the study by Belbase et al. Belbase et al. (2023), we observe 2 bands spanning a larger portion of the Brillouin zone, i.e. the band at -0.6 eV along -M-K- and the band at -0.8 eV along A-L-H-A. The orbital-projected spectral functions, reported in the SM sup , allow us to identify the dominant character of these bands. The local projectors are constructed by real spherical harmonics defined on a rotated reference frame where lays in the kagome plane and points toward the neighboring Sn atom, as illustrated in Figure S3 of the SM sup . This choice is analogous to the one used for Sc3Mn3Al7Si5 in a very recent DFT+DMFT study Samanta et al. (2024). With the help of these data, we can establish that the flat band at -0.6 eV has mostly | majority-spin character, which is consistent with recent results on the related breathing kagome material Ni3In Ye et al. (2024). This can be easily verified through a further rotation of the local orbitals to match the two reference frames. The other flat band, at -0.8 eV, exhibits a similar character, but has also a substantial contribution from | majority-spin states, in particular between L and H. Along the same high-symmetry directions, one can also observe the corresponding minority spin bands at higher energies, namely 1.6 eV and 1.4 eV, see Figure S10 of the SM sup . Another notable flat band can be observed at -1.2 eV, mainly of | minority-spin character. Due to the strong interlayer coupling between the two kagome layers Lin et al. (2018); Xu et al. , these flat bands do not extend over the entire Brillouin zone, but acquire a strong dispersion along the -A direction. In Figure 4(a), we also observe another set of non-dispersive bands extending along smaller segments of the Brillouin zone, mainly at smaller excitation energies. The two bands highlighted in blue at about 50 meV and 200 meV come from the states determining the nodal plane between the U∗ points, which was suggested to drive a large anomalous Nersnt effect, ten times larger than for pure Fe Chen et al. (2022). For these bands the hybridization between the Fe- states and Sn- states becomes more important, as it can be inferred from comparing the corresponding orbital-projected spectral functions in Figures S4 and S7 of the SM sup . In general, the presence of flat bands has been linked to the formation of ferromagnetic order in Fe stannites, especially for the analysis of Fe3Sn2, whose kagome features are enhanced by its quasi-two-dimensional nature Tasaki (2008); Lin et al. (2018); Pollmann et al. (2008). In Fe3Sn, the role played by the flat bands and other non-dispersive states in the long-range order is expected to be less crucial, due to the three-dimensional magnetism emphasized above. Moreover, the analysis of Figure 4(a) as well as the orbital projections reported in the SM sup has shown that the carriers mediating the exchange coupling belong to hybridized - states having a marked itinerant character, at least in DFT.
@@
The inclusion of SOC does not alter the shape and position of the non-dispersive bands significantly, but results in gap opening at the band-crossing points, which we highlighted in orange color in Figure 4(d). Among the opened gaps, one can find a number of Weyl nodes, as recently demonstrated by Belbase et al. Belbase et al. (2023). These points were suggested to be the cause of the large anomalous Hall conductivity observed in experiment Shen et al. (2022), but electronic structure calculations indicate more structured contributions across the entire Brillouin zone Belbase et al. (2023). By changing the magnetization axis to the hard axis (001), we observe a notable band splitting close to the Fermi level, as shown in Figures S8 and S9 of the SM sup . This splitting stems from the dependence of the electron scattering on the spin orientation and differs from the Zeeman splitting. This phenomenon is recognized as the giant magneto band structure effect, as discussed in literature for two-dimensional CrI3 Jiang et al. (2018). Furthermore, for a magnetization oriented along (001), one can observe a different set of Weyl nodes, regarding both energy and distribution in the Brillouin zone Belbase et al. (2023); Fu et al. (2025).
@@
Next, we focus on the band structure (spectral function) obtained in DFT+DMFT, shown in panels (b) and (e) of Figure 4. Due to the imaginary part of the self-energy, the quasiparticles acquire a finite lifetime which makes the energy bands well-defined only for energies close to the Fermi level, consistently with the Fermi liquid theory Grechnev et al. (2007). This is also evident from the PDOS shown in Figure 3, where no structure can be resolved below -2 eV and above +2 eV. The real part of the self-energy renormalizes the DFT eigenvalues, which is particularly evident in Figure 4(b) for the non-dispersive bands highlighted in blue, which move much closer to the Fermi level. A similar renormalization has been reported recently for other kagome metals, such as Mn3Sn Yu et al. (2022) and AV3Sb5 (A = K, Cs, Rb) Hasan et al. (2023). The proper flat bands, highlighted in red, also move closer to the Fermi level. All these shifts are due to the combination of Fermi liquid renormalization, which induces a band narrowing Kuneš et al. (2009), and a rigid shift of the Fe- orbitals, driven by a larger orbital polarization and a decreased spin splitting, as emphasized in the Section III. To compare these two effects, we can inspect the results obtained in DFT+U, shown in panels (c) and (f) of Figure 4. Since these plots do not contain the effect of Fermi liquid renormalization, we can conclude that the latter is quite substantial in the correlated electronic structure obtained from DFT+DMFT, at least in reference to the energy scale we are focusing on. By extracting the quasiparticle weight from the self energy, we obtain a mass renormalization of 1.31 for majority spin and 1.25 for minority spin. These values are in line with those obtained for ferromagnetic kagome metals such as AV3Sb5 (A = K, Cs, Rb) Liu et al. (2022); Hasan et al. (2023); Zhao et al. (2021b) and ScV6Sn6 Yu et al. (2024), but are slightly smaller than those expected for Mn3Sn Samanta et al. (2024); Yu et al. (2022). This is consistent with previous findings that the magnetic fluctuations in Mn tend to be larger than in Fe, as the former is closer to having a half-filled shell Di Marco et al. (2009); Di Marco et al. (2009). For completeness, the orbital-decomposed renormalization is reported in Table 3.
@@
Interestingly, for both DFT+DMFT and DFT+U, the band segments along -M and M-K are predicted to cross the Fermi level, while those along A-L move at very small excitation energies. However, we should stress that this effect is very sensitive to the precise choice of the Coulomb interaction parameters. For example, the DFT+DMFT results for values of and that provide the optimal MAE, discussed in Section III, do not exhibit these changes (data not shown). This uncertainty, as well as having neglected defects of any sort, makes it impossible to provide a quantitative analysis of how far these non-dispersive bands are going to be located from the Fermi level. Nevertheless, our results highlight the following fundamental aspects: 1) a series of non-dispersive bands are located in the very vicinity of the Fermi level; 2) these non-dispersive bands are strongly affected by the local Coulomb repulsion. This suggests that one could induce a variety of Lifshitz transitions by a small modulation of external parameters, which could take the form of a small strain (to change the bandwidth) or doping charge carriers (to change the position of the Fermi level). We believe that these aspects are behind the large sensitivity to doping and growth conditions exhibited by the magnetic properties, which have prevented the stabilization of the MAE along the out-of-plane direction Fayyazi et al. (2019). Interestingly, the formation of unexpected hole pockets and band crossing analogous to the one observed here has recently been reported by laser-based micro-focused ARPES measurements of Fe3Sn2 Ekahana et al. (2024).
Finally, we focus on the spectral functions obtained in DFT+DMFT and DFT+U when SOC is included, as observed in Figure 4(e) and Figure 4(f) respectively. We can see that the Weyl point close to the M point moves in a way to open a gap precisely at the Fermi level, while the one just close to it moves to lower energies. This new configuration may alter the contributions of these points to the Berry curvature, thus bringing a quantitative change to the topological properties. This may provide another route to explain the remaining discrepancy reported between calculated and theoretical values for the anomalous Hall conductivity in both single crystals Low et al. (2023); Belbase et al. (2023) and polycrystalline samples Chen et al. (2022); Fu et al. (2025). The changes induced in the interlayer coupling, visible through an additional crossing point along the -A direction with respect to DFT are likely behind the drastic increase of the Dzyaloshinskii-Moriya interaction, reported in Figure 2(b).

### v2

@@ III Magnetic properties
The amplitude of the DM vector as a function of the interatomic distance is shown in Figure 2(b), while the components Dx, Dy, and Dz for selected exchange paths are given in Table 2. The largest coupling is interlayer, connecting Fe atoms along the path E2, and amounts to 0.52 meV, mostly arising from its Dz component. Inside the kagome plane, the first and second shells of NNs are characterized by slightly smaller DM vectors, whose major contributions still come from the Dz component. These finite values of the DM interaction indicate the tendency for chirality between two Fe sites with the preferred direction for the DM vector along the z-axis (out-of-plane). The magnitude of the DM interaction is relatively small (3-4%) compared to the isotropic Heisenberg exchange tensor, making the likelihood of spin canting quite low [75]. Definitive confirmation of collinear order would require the solution of the effective Heisenberg Hamiltonian defined by Eqn. 2, which would be an interesting task for future works involving multi-scale modeling [76]. Furthermore, we see that |D| for the forth and fifth NN shells is zero, because of the presence of inversion symmetry, as evidenced by the interlayer exchange paths E4 and E5 depicted in Figure 1(c).
DFT | DFT+DMFT | DFT+U | ||||
| E1 | Dx | 0 | 0 | 0 | ||||
| Dy | 0 | 0 | 0 | |||||
| Dz | 0.41 | 0.22 | 0.50 | |||||
| E2 | Dx | -0.08 | -0.05 | -0.05 | ||||
| Dy | 0.04 | 0.02 | 0.02 | |||||
| Dz | 0.51 | 0.19 | 0.36 | |||||
| E3 | Dx | 0 | 0 | 0 | ||||
| Dy | 0 | 0 | 0 | |||||
| Dz | 0.28 | 0.56 | 0.78 | |||||
| E6 | Dx | -0.16 | -0.27 | -0.31 | ||||
| Dy | 0 | 0 | 0 | |||||
| Dz | 0 | 0 | 0 | |||||
| E9 | Dx | 0.09 | 0.09 | 0.13 | ||||
| Dy | 0.16 | 0.17 | 0.23 | |||||
@@
The inclusion of an explicit Hubbard term via DFT+U and DFT+DMFT leads to a small reduction of the isotropic exchange, as shown in Figure 2(a). This reduction reflects the smaller values of the Fe spin moments reported in Table 1 through the definition used in Eqn. 2. By inspecting the projected density of states (PDOS) for the Fe- states, depicted in Figure 3, we observe that the exchange splitting is largest in DFT and smallest in DFT+DMFT, in agreement with the trends of the magnetic moments and interatomic interactions. The overall reduction of the strength of the magnetic coupling in DFT+DMFT is similar to the one previously observed for bcc Fe [77, 57]. Nevertheless, it is surprising that the DFT+U does not cause a stronger localization of the Fe- electrons, making the ’s shorter ranged. Although a possible reason for the observed behavior may be the usage of smaller Coulomb interaction parameters in DFT+U, additional calculations demonstrate that this is not the leading factor (data not shown). In fact, the long-range nature is found to be a consequence of the AMF double counting, which is usually used for correlated metals [78]. To evaluate the impact of the double counting, we performed additional electronic structure calculations in DFT+U in the fully-localized limit (FLL) while keeping the other settings unchanged. As expected, a much stronger electronic localization is observed, resulting in an increase of the exchange splitting (data not shown), a higher local magnetic moment of 2.54 B, and an increase of the first NN coupling up to 31 meV. This magnetic moment deviates significantly from the experimentally reported values, which demonstrates that the FLL double counting is not suitable for describing this system, as usual for itinerant ferromagnets [78].
@@
The inclusion of the on-site Coulomb interaction via DFT+DMFT and DFT+U leads to significant qualitative and quantitative changes in the DM interaction. For both methods, Figure 2(b) shows a marked increase in the value of |D| for the third shell of NNs, corresponding to the intralayer exchange path E3. The first shell of NNs, along the intralayer exchange path E1, exhibits a more complex trend, as the DM interaction increases in DFT+U but decreases in DFT+DMFT. The dominant character of the intralayer terms remains Dz, compatibly with the symmetry of the kagome lattice. The DM interaction with the second shell of NNs, along the exchange path E2, decreases for both DFT+U and DFT+DMFT, an effect driven mainly by a change in Dz. Notably, there is also a significant increase in the magnitude of the DM interaction with farther NNs, as e.g. in the sixth and ninth shells, which we attribute to a stronger orbital polarization driven by the inclusion of the on-site Coulomb interaction. These data, and in particular the decomposition shown in Table 2, suggest deeper changes in the electronic structure than those observed from the PDOS of Figure 3. We will investigate these changes in the next section.
@@
For completeness, we also calculated the values of the interatomic exchange interactions for the non-relativistic case, corresponding to Eqn. 1. As illustrated in Figure S1 of the SM [66], the isotropic exchange is only marginally affected by the presence of SOC, and the primary consequence of the latter is to induce anisotropic terms as the DM interaction, favoring spin canting and complex magnetic textures [7].
@@ IV Details of the electronic structure
The spectral functions along high-symmetry directions in the Brillouin zone are reported in Figure 4 for DFT, DFT+DMFT and DFT+U, with and without the inclusion of SOC. As in the previous Section, the data for SOC are shown for a magnetization along the easy axis (100). Focusing on the DFT results, panels (a) and (d), we see the presence of several bands with a relatively flat dispersion near the Fermi level. For convenience, we highlighted those bands in magenta and blue colors, reflecting their spin character. In agreement with the study by Belbase et al. [3], we observe 2 bands spanning a larger portion of the Brillouin zone, i.e. the band at -0.6 eV along -M-K- and the band at -0.8 eV along A-L-H-A. The orbital-projected spectral functions, reported in the SM [66], allow us to identify the dominant character of these bands. The local projectors are constructed by real spherical harmonics defined on a rotated reference frame where lays in the kagome plane and points toward the neighboring Sn atom, as illustrated in Figure S3 of the SM [66]. This choice is analogous to the one used for Sc3Mn3Al7Si5 in a very recent DFT+DMFT study [29]. With the help of these data, we can establish that the flat band at -0.6 eV has mostly | majority-spin character, which is consistent with recent results on the related breathing kagome material Ni3In [79]. This can be easily verified through a further rotation of the local orbitals to match the two reference frames. The other flat band, at -0.8 eV, exhibits a similar character, but has also a substantial contribution from | majority-spin states, in particular between L and H. Along the same high-symmetry directions, one can also observe the corresponding minority spin bands at higher energies, namely 1.6 eV and 1.4 eV, see Figure S10 of the SM [66]. Another notable flat band can be observed at -1.2 eV, mainly of | minority-spin character. Due to the strong interlayer coupling between the two kagome layers [13, 80], these flat bands do not extend over the entire Brillouin zone, but acquire a strong dispersion along the -A direction. In Figure 4(a), we also observe another set of non-dispersive bands extending along smaller segments of the Brillouin zone, mainly at smaller excitation energies. The two bands highlighted in blue at about 50 meV and 200 meV come from the states determining the nodal plane between the U∗ points, which was suggested to drive a large anomalous Nernst effect, ten times larger than for pure Fe [15]. For these bands the hybridization between the Fe- states and Sn- states becomes more important, as it can be inferred from comparing the corresponding orbital-projected spectral functions in Figures S4 and S7 of the SM [66]. In general, the presence of flat bands has been linked to the formation of ferromagnetic order in Fe stannites, especially for the analysis of Fe3Sn2, whose kagome features are enhanced by its quasi-two-dimensional nature [19, 13, 81]. In Fe3Sn, the role played by the flat bands and other non-dispersive states in the long-range order is expected to be less crucial, due to the three-dimensional magnetism emphasized above. Moreover, the analysis of Figure 4(a) as well as the orbital projections reported in the SM [66] has shown that the carriers mediating the exchange coupling belong to hybridized - states having a marked itinerant character, at least in DFT.
@@
The inclusion of SOC does not alter the shape and position of the non-dispersive bands significantly, but results in gap opening at the band-crossing points, which we highlighted in orange color in Figure 4(d). This means that the Weyl nodes move away from the high-symmetry directions, as recently demonstrated by Belbase et al. [3]. These points were suggested to be the cause of the large anomalous Hall conductivity observed in experiment [69], but electronic structure calculations indicate more structured contributions across the entire Brillouin zone [3]. By changing the magnetization axis to the hard axis (001), we observe a notable band splitting close to the Fermi level, as shown in Figures S8 and S9 of the SM [66]. This splitting stems from the dependence of the electron scattering on the spin orientation and differs from the Zeeman splitting. This phenomenon is recognized as the giant magneto band structure effect, as discussed in literature for two-dimensional CrI3 [82]. Furthermore, we observe that the distribution of the gap openings depends significantly on the magnetization direction. This difference is connected to a different shift of the Weyl nodes under SOC, regarding both energy and location in the Brillouin zone [3, 83].
@@
Next, we focus on the band structure (spectral function) obtained in DFT+DMFT, shown in panels (b) and (e) of Figure 4. Due to the imaginary part of the self-energy, the quasiparticles acquire a finite lifetime which makes the energy bands well-defined only for energies close to the Fermi level, consistently with the Fermi liquid theory [41]. This is also evident from the PDOS shown in Figure 3, where no structure can be resolved below -2 eV and above +2 eV. The real part of the self-energy renormalizes the DFT eigenvalues, which is particularly evident in Figure 4(b) for the non-dispersive bands highlighted in blue, which move much closer to the Fermi level. A similar renormalization has been reported recently for other kagome metals, such as Mn3Sn [84] and AV3Sb5 (A = K, Cs, Rb) [6]. The proper flat bands, highlighted in red, also move closer to the Fermi level. All these shifts are due to the combination of Fermi liquid renormalization, which induces a band narrowing [85], and a rigid shift of the Fe- orbitals, driven by a larger orbital polarization and a decreased spin splitting, as emphasized in the Section III. To compare these two effects, we can inspect the results obtained in DFT+U, shown in panels (c) and (f) of Figure 4. Since these plots do not contain the effect of Fermi liquid renormalization, we can conclude that the latter is quite substantial in the correlated electronic structure obtained from DFT+DMFT, at least in reference to the energy scale we are focusing on. By extracting the quasiparticle weight from the self energy, we obtain a mass renormalization of 1.31 for majority spin and 1.25 for minority spin. These values are in line with those obtained for ferromagnetic kagome metals such as AV3Sb5 (A = K, Cs, Rb) [86, 6, 87] and ScV6Sn6 [88], but are slightly smaller than those expected for Mn3Sn [29, 84]. This is consistent with previous findings that the magnetic fluctuations in Mn tend to be larger than in Fe, as the former is closer to having a half-filled shell [47, 43]. For completeness, the orbital-decomposed renormalization is reported in Table 3.
@@
Interestingly, for both DFT+DMFT and DFT+U, the band segments along -M and M-K are predicted to cross the Fermi level, while those along A-L move at very small excitation energies. However, we should stress that this effect is very sensitive to the precise choice of the Coulomb interaction parameters. For example, the DFT+DMFT results for values of and that provide the optimal MAE, discussed in Section III, do not exhibit these changes (data not shown). This uncertainty, as well as having neglected defects of any sort, makes it impossible to provide a quantitative analysis of how far these non-dispersive bands are going to be located from the Fermi level. Nevertheless, our results highlight the following fundamental aspects: 1) a series of non-dispersive bands are located in the very vicinity of the Fermi level; 2) these non-dispersive bands are strongly affected by the local Coulomb repulsion. This suggests that one could induce a variety of Lifshitz transitions by a small modulation of external parameters, which could take the form of a small strain (to change the bandwidth) or doping charge carriers (to change the position of the Fermi level). We believe that these aspects are behind the large sensitivity to doping and growth conditions exhibited by the magnetic properties, which have prevented the stabilization of the MAE along the out-of-plane direction [23]. Interestingly, the formation of unexpected hole pockets and band crossing analogous to the one observed here has recently been reported by laser-based micro-focused ARPES measurements of Fe3Sn2 [30].
Finally, we focus on the spectral functions obtained in DFT+DMFT and DFT+U when SOC is included, as observed in Figure 4(e) and Figure 4(f) respectively. We can see that the band crossing closest to the M point, corresponding to a Weyl point in the analysis by Belbase et al. [3], moves off the high-symmetry direction, leading to the opening of a gap precisely at the Fermi level. Conversely, the band crossing close to it just moves to lower energies (-0.2 eV for both DFT+DMFT and DFT+U). This new configuration may alter the contributions of these points to the Berry curvature, thus bringing a quantitative change to the topological properties. This may provide another route to explain the remaining discrepancy reported between calculated and theoretical values for the anomalous Hall conductivity in both single crystals [26, 3] and polycrystalline samples [15, 83]. The changes induced in the interlayer coupling, visible through an additional crossing point along the -A direction with respect to DFT are likely behind the drastic increase of the Dzyaloshinskii-Moriya interaction, reported in Figure 2(b).